# Differentially Private Matchings

### v2

@@ A.1 Concentration Inequalities
Given a random variable $X\sim Lap(b)$ drawn from a Laplace distribution with expectation 00, the probability $|X|>c\ln(n)$ is $n^{-\frac{c}{b}}$.
@@
Let $X=\sum_{i=1}^{n}X_{i}$ where each $X_{i}$ is a Bernoulli variable which takes value $1$ with probability $p_{i}$ and 00 with probability $1-p_{i}$. Let $\mu=\mathbb{E}[X]=\sum_{i=1}^{n}p_{i}$. Then, it holds:
@@
    Lower Tail: $\mathsf{Pr}(X\leq(1-\psi)\cdot\mu)\leq\exp\left(-\frac{\psi^{2}\mu}{3}\right)$ for all $10<\psi<$ 1.
@@
We now restate and prove Lemma 2.14. See 2.14
@@ A.3 Multi-Response Sparse Vector Technique
[…]
We use the class $\textsc{SVT}(\varepsilon,\Delta,c)$ (Algorithm 10) where $\varepsilon$ is our privacy parameter, $\Delta$ is an upper bound on the maximum sensitivity of incoming queries, and $c$ is the maximum number of “above” queries we can make. The class provides a $\textsc{ProcessQuery}(query,threshold)$ function where $query$ is the query to SVT and $threshold$ is the threshold that we wish to check whether the query exceeds.
@@
1 Input: privacy budget $\varepsilon$, upper bound on query sensitivity $\Delta$, maximum allowed “above” answers $c$
@@ 13                   $\operatorname{count}\leftarrow\operatorname{count}+1$
Appendix B Lower Bound for Explicit Fractional Matchings
We now prove Theorem 1.2 which is restated below for convenience: in the basic 0/1 weight $(\varepsilon,\delta)$-edge DP model (Definition 2.7), the required error is essentially linear, even if we allow for the algorithm to output a fractional $b$-matching for very large $b$. This proof is essentially the same as the proof of Theorem 1.1, but we give it explicitly for completeness.
See 1.2
Given a graph $G=(V,E)$, let $\nu^{f}(G)$ denote the size of the largest fractional matching in $G$ Given a (randomized) algorithm $\mathcal{A}$ which outputs some fractional $b$-matching $X_{G}^{\mathcal{A}}$ when run on input graph $G=(V,E)$, i.e. $X_{G}^{\mathcal{A}}\in[0,1]^{\binom{V}{2}}$ is a random vector satisfying $\sum_{v\neq u}X_{G,uv}^{\mathcal{A}}\leq b$ for all $u\in V$ with probability 1, let $\nu_{\mathcal{A}}^{f}(G)=\nu^{f}((V,X_{G}^{\mathcal{A}}\big{|}_{E}))$ be the size of a maximum fractional matching in the graph restricted to edges that are in $G$.
We now begin our proof of Theorem 1.2. At a very high level, we will give a collection of input graphs and argue that any differentially private algorithm must be a poor approximation on at least one of them. This is in some ways similar to the standard “packing arguments” used to prove lower bounds for differentially private algorithms (see, e.g., [DMN23]), but instead of using a packing / averaging argument we will instead use a symmetry argument. We will argue that without loss of generality, any DP algorithm for our class of inputs must be highly symmetric in that $\text{E}[X_{G,uv}^{\mathcal{A}}]$ for each non-edge $uv\notin E$ is identical. This makes it easy to argue that high-quality algorithms cannot satisfy differential privacy.
We begin with a description of our inputs and some more notation. Given $n$ nodes $V$ for an even integer $n>0$, let $\mathcal{M}$ denote the set of all perfect matchings on $V$. Let $\mathcal{A}$ be an $(\varepsilon,\delta)$-differentially private algorithm which always returns a fractional $b$-matching $X_{G}^{\mathcal{A}}$ on $V$ as defined above. For $G\in\mathcal{M}$ and $uv\in\binom{V}{2}$, let $\text{E}[X_{G,uv}^{\mathcal{A}}]$ denote the expected fractional value that $\mathcal{A}(G)$ assigns to $uv$. Let $\nu_{\mathcal{A}}^{f}=\min_{G\in\mathcal{M}}\text{E}[\nu_{\mathcal{A}}^{f}(G)]$ denote the worst-case expected utility of $\mathcal{A}$ over $\mathcal{M}$. Note that for $G\in\mathcal{M}$, every subset of $E(G)$ is a matching. By linearity of expectations, this is equivalent to saying that $\sum_{uv\in E(G)}{\text{E}[X_{G,uv}^{\mathcal{A}}]}\geq\nu_{\mathcal{A}}^{f}$ for all $G\in\mathcal{M}$.
Lemma B.1.
If $\mathcal{A}$ satisfies $(\varepsilon,\delta)$-DP and has expected utility at least $\nu_{\mathcal{A}}^{f}$ over $\mathcal{M}$, then there is an algorithm $\mathcal{A}^{\prime}$ which also satisfies $(\varepsilon,\delta)$-DP and has $\nu_{\mathcal{A}^{\prime}}^{f}\geq\nu_{\mathcal{A}}^{f}$. Moreover $\mathcal{A}^{\prime}$ satisfies the following symmetry property: $\text{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]=\text{E}[X_{G,ab}^{\mathcal{A}^{\prime}}]$ for all $uv,ab\not\in E(G)$.
Proof.
Consider the following algorithm $\mathcal{A}^{\prime}$. We first choose a permutation $\pi$ of $V$ uniformly at random from the space $S_{V}$ of all permutations of $V$. Let $\pi(G)$ denote the graph isomorphic to $G$ obtained by applying this permutation, i.e., by creating an edge $\{\pi(u),\pi(v)\}$ for each edge $\{u,v\}\in E(G)$. We run $\mathcal{A}$ on $\pi(G)$, and let $\pi(H)$ denote the resulting fractional $b$-matching. We then “undo” $\pi$ to get the fractional $b$-matching $X_{G,uv}^{\mathcal{A}^{\prime}}=X_{\pi(G),\pi(u)\pi(v)}^{\mathcal{A}}$. We return $X_{G}^{\mathcal{A}^{\prime}}$.
We first claim that $\mathcal{A}^{\prime}$ is $(\varepsilon,\delta)$-DP. This is straightforward, since it simply runs $\mathcal{A}$ on a uniformly random permuted version of $G$, and then does some post-processing. Slightly more formally, for any neighboring graphs $G$ and $G^{\prime}$ and for any $\pi\in S_{V}$, the graphs $\pi(G)$ and $\pi(G^{\prime})$ are also neighboring graphs. Hence the differential privacy guarantee of $\mathcal{A}$ implies that running $\mathcal{A}$ on $\pi(G)$ and $\pi(G^{\prime})$ is $(\varepsilon,\delta)$-DP, and then “undoing” $\pi$ is simply post-processing. So $\mathcal{A}^{\prime}$ has the same privacy guarantees as $\mathcal{A}$.
$\text{E}[\nu_{\mathcal{A}^{\prime}}^{f}(G)]$ $=\frac{1}{n!}\sum_{\pi\in S_{V}}\text{E}[\nu_{\mathcal{A}}^{f}(\pi(G))]\geq\frac{1}{n!}\sum_{\pi\in S_{V}}\nu_{\mathcal{A}}^{f}=\nu_{\mathcal{A}}^{f}\,.$
Since this is true for all $G\in\mathcal{M}$, we have that $\nu_{\mathcal{A}^{\prime}}=\min_{G\in\mathcal{M}}\text{E}[\nu_{\mathcal{A}^{\prime}}(G)]\geq\nu_{\mathcal{A}}$, as claimed.
We now prove the symmetry property. Note that there are $n!$ permutations in $S_{V}$, but only $\frac{n!}{2^{n/2}(n/2)!}$ different perfect matchings (i.e., $|\mathcal{M}|=\frac{n!}{2^{n/2}(n/2)!}$). This is because $\pi(G)$ can equal $\pi^{\prime}(G)$ for many for many $\pi\neq\pi^{\prime}$. In particular, it is not hard to see that |$|\{\pi\in S_{V}:\pi(G)=G^{\prime}\}|=2^{n/2}(n/2)!$ for all $G,G^{\prime}\in\mathcal{M}$. Fix $G,G^{\prime}\in\mathcal{M}$, and let $\Pi(G,G^{\prime})=\{\pi\in S_{V}:\pi(G)=G^{\prime}\}$. It is not hard to see that if we draw a permutation $\pi$ uniformly at random from $\Pi(G,G^{\prime})$, then for any $\{u,v\}\not\in E(G)$, the pair $\{\pi(u),\pi(v)\}$ is uniformly distributed among the set $\{\{u^{\prime},v^{\prime}\}\not\in E(G^{\prime})\}$. Thus we have that
$\text{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]$ $=\frac{1}{n!}\sum_{\pi\in S_{V}}\text{E}[X_{\pi(G),\pi(u)\pi(v)}^{\mathcal{A}}]$
$=\frac{1}{n!}\sum_{G^{\prime}\in\mathcal{M}}\sum_{\pi\in\Pi(G,G^{\prime})}\text{E}[X_{G^{\prime},\pi(u)\pi(v)}^{\mathcal{A}}]$
$=\frac{1}{n!}\sum_{G^{\prime}\in\mathcal{M}}\sum_{ab\not\in E(G^{\prime})}\text{E}[X_{G^{\prime},ab}^{\mathcal{A}}].$
Hence $\text{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]$ is actually independent of $u,v$, and thus for all $uv,ab\not\in E(G)$, we have that $\text{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]=\text{E}[X_{G,ab}^{\mathcal{A}^{\prime}}]$ as claimed. ∎
Lemma B.1 implies that if we can prove an upper bound on the expected utility for DP algorithms that obey the symmetry property, then the same upper bound applies to all DP algorithms. We will now prove such a bound.
Lemma B.2.
Let $\mathcal{A}$ be an algorithm which is $(\varepsilon,\delta)$-DP, always outputs a fractional $b$-matching, and satisfies the symmetry property of Lemma B.1. Then
$\nu_{\mathcal{A}}^{f}\leq e^{2\varepsilon}b+\delta n.$
Proof.
Fix some graph $G\in\mathcal{M}$, and let $uv,ab\in E(G)$. Let $G^{\prime}$ be the graph obtained by removing $uv$ and $ab$, and adding $ub$ and $av$. Note that $G$ and $G^{\prime}$ are distance $2$ away from each other.
For all $y\neq u,b$, note that $uy\notin E(G^{\prime})$. Thus it must be the case that $\text{E}[X_{G,uy}^{\mathcal{A}}]\leq b/(n-2)$, since otherwise the symmetry property of Lemma B.1 and linearity of expectations implies that the expected (fractional) degree of $u$ in the output of $\mathcal{A}(G)$ is at least $\sum_{y\neq u,b}\text{E}[X_{G,uy}^{\mathcal{A}}]>(n-2)\frac{b}{n-2}=b$, contradicting our assumption that $\mathcal{A}$ always outputs a fractional $b$-matching. Thus $\text{E}[X_{G^{\prime},uv}^{\mathcal{A}}]\leq b/(n-2)$. Since $G$ and $G^{\prime}$ are at distance $2$ from each other, this implies that
$\text{E}[X_{G,uv}^{\mathcal{A}}]$ $=\int_{0}^{1}x\cdot\Pr[X_{G,uv}^{\mathcal{A}}=x]\mathop{}\!\mathrm{d}x$
$\leq\int_{0}^{1}x\left(e^{2\varepsilon}\cdot\Pr[X_{G^{\prime},uv}^{\mathcal{A}}=x]+2\delta\right)\mathop{}\!\mathrm{d}x$
$=e^{2\varepsilon}\cdot\text{E}[X_{G^{\prime},uv}^{\mathcal{A}}]+\delta$
$\leq e^{2\varepsilon}\frac{b}{n-2}+\delta\,.$
Since this argument applies to all $uv\in E(G)$, we have that
$\text{E}[\nu_{\mathcal{A}}^{f}(G)]$ $=\sum_{uv\in E(G)}\text{E}[X_{G,uv}^{\mathcal{A}}]$
$\leq\sum_{uv\in E(G)}\left(e^{2\varepsilon}\frac{b}{n-2}+\delta\right)$
$\leq e^{2\varepsilon}\frac{bn}{2(n-2)}+\delta\frac{n}{2}.$
$\leq e^{2\varepsilon}b+\delta n.$
Since $\nu_{\mathcal{A}}^{f}=\min_{G\in\mathcal{M}}\text{E}[\nu_{\mathcal{A}}^{f}(G)]$, this implies that $\nu_{\mathcal{A}}^{f}\leq e^{2\varepsilon}b+\delta n$, as claimed. ∎
We can now use these two lemmas to prove Theorem 1.2.
Proof of Theorem 1.2.
The combination of Lemma B.1 and Lemma B.2 imply that any $(\varepsilon,\delta)$-DP algorithm $\mathcal{A}$ which always outputs a fractional $b$-matching has $\nu_{\mathcal{A}}^{f}\leq e^{2\varepsilon}b+\delta n$. We also know that $\nu^{f}(G)=n/2$ for all $G\in\mathcal{M}$. Thus by the definition of an $(\alpha,\beta)$-approximation, it must be the case that $\alpha(e^{2\varepsilon}b+\delta n)+\beta\geq n/2$. ∎

### v3

@@ A.1 Concentration Inequalities
Given a random variable $X\sim Lap(b)$ drawn from a Laplace distribution with expectation 0, the probability $|X|>c\ln(n)$ is $n^{-\frac{c}{b}}$.
@@
Let $X=\sum_{i=1}^{n}X_{i}$ where each $X_{i}$ is a Bernoulli variable which takes value $1$ with probability $p_{i}$ and 0 with probability $1-p_{i}$. Let $\mu=\mathbb{E}[X]=\sum_{i=1}^{n}p_{i}$. Then, it holds:
@@
    Lower Tail: $\mathsf{Pr}(X\leq(1-\psi)\cdot\mu)\leq\exp\left(-\frac{\psi^{2}\mu}{3}\right)$ for all $0<\psi<1$.
@@
We now restate and prove Lemma 3.15. See 3.15
@@ A.3 Multi-Response Sparse Vector Technique
We recall a variant of the sparse vector technique that enables multiple responses as introduced by [lyu2017understanding]. Let $D$ be an arbitrary (graph) dataset, $(f_{t},\tau_{t})$ a sequence of (possibly adaptive) query-threshold pairs, $\Delta$ an upper bound on the maximum sensitivity of all queries $f_{t}$, and an upper bound $c$ on the maximum number of queries to be answered “above”. Typically, the AboveThreshold algorithm stops running at the first instance of the input exceeding the threshold, but we use the variant where the input can exceed the threshold at most $c$ times where $c$ is a parameter passed into the function.
We use the class $\textsc{SVT}(\varepsilon,\Delta,c)$ (Algorithm 9) where $\varepsilon$ is our privacy parameter, $\Delta$ is an upper bound on the maximum sensitivity of incoming queries, and $c$ is the maximum number of “above” queries we can make. The class provides a $\textsc{ProcessQuery}(query,threshold)$ function where queryquery is the query to SVT and thresholdthreshold is the threshold that we wish to check whether the query exceeds.
@@
1Input: privacy budget $\varepsilon$, upper bound on query sensitivity $\Delta$, maximum allowed “above” answers $c$